Nadyah:

If f(x-1)= 2x+3 for all values of x, what is the value of f(-3)

KpopAsianDramaFanYT:

Substitute -3 for x

Florisalreadytaken:

the reply above is just complicated for no reason. so, if \(f(x-1)= 2x+3\) for all values of x, what is the value of\( f(-3)\) well we want to get that \( f(x-1) \) into \( f(-3) \) to do that, we have to put in \( -2\) for \(x\), so that it is true. from here, its just simple math: \( f(-2-1) = 2 \times -2+3 \) \( f(-3)=-4+3 \) \( f(-3)=-1 \)